The length of a rectangle is three times the width. If the perimeter is 36
cm, find the width.

Answer by ewatrrr(24785)

 
Hi,
.
L = 3w, P = 36cm
2L + 2w = P
6w + 2w = 36cm
8w = 36cm
w = 4.5cm
L = 13.5cm
Checking
2L + 2w = P
27cm + 9cm = 36cm CHECKS!